Meeting:2011/05/12

From SOWNWiki
Jump to: navigation, search

Meeting (to be) held on 2011/05/12 at 18:00 in Zepler CLS Lecture Room

Previous meeting: 5 May 2011 18:30:00 Next meeting: 19 May 2011 18:00:00


Agenda

Amateur Radio (18:00-18:45)



  • AOB

Joint Agenda (18:45-19:15)




  • AOB


SOWN (19:15-20:00)



  • Icinga Alerts (Crwilliams)
    • Monitor/Debsums
    • Dev/Debsums
    • VPN/Ping
    • UPS checks and Auth/UPS
    • SUSU/Boottime-SNMP





  • AOB

Minutes

Amateur Radio

  • FT101 Status (HeliosFA)
    • Currently broken, due to stuck pot.
    • (HeliosFA) fix in progress
  • Members (HeliosFA)
    • No new members
    • But a few people showed interest
  • Club outing (HeliosFA)
    • Possible outdoor trip to more radio-friendly location
      • Milton-on-sea
      • Wayfarer's Walk
      • Sea-side locations
      • High locations
  • Antenna Building Evening (HeliosFA)
    • At SOWN workshop
    • Or over next few SUWS meetings
  • AOB
    • Club callsign
      • Ofcom contacted, email sent to correct station address.

Joint Agenda

  • SUWS Website (HeliosFA)
    • Colour scheme is too dark, contrast too low
    • Link colouring needs fixing
    • Light grey good
    • Check of server hosting rules, inc. DNS
  • Membership Fees Vote (Davetaz)
    • No redistibution
    • Motion Carried
    • Student fees include postgraduate students
  • RSGB Affiliation (HeliosFA)
    • Affiliation vote carried
  • AOB

SOWN (19:15-20:00)

  • May Festival Review (daveruss)
    • Sort area for Fresher's week
    • Could do a repeat of May Fair setup
    • Justify inside/outside setup as one stand, one week long
  • Admin Tasks
    • Issues changing SSID of 8021X interface (daveruss)
    • Cron emails (Leth)
      • (Leth) still working on it
      • partially complete
      • check mailscanner ipv6
    • Undercroft situation (Crwilliams)
      • Last kit moved out.
      • Cupboard now locked with another lock
      • Inventory not updated
      • Scorpia to chase up SUSU for storage via email with Crwilliams
    • Labelling Inventory (Davetaz)
    • Kit order (Davetaz)
      • no progress
    • DNS record usage by servers (Davetaz)
      • routing based on source address will not resolve problem
      • removed problematic route on monitor
      • workshop task: look at ipv6
  • HomeNodes (HeliosFA)
    • Stags Head - reported issues with 802.1X
      • Workshop task: look at during lunch break.
    • Chamberlain
      • No progress, HeliosFA still waiting for reply.
      • Get node back.
  • Icinga Alerts (Crwilliams)
    • Monitor/Debsums
      • Checks passive and fixed
    • Dev/Debsums
      • Checks passive and fixed
    • VPN/Ping
      • Monitor had incorrect route to itself, route removed
    • UPS checks and Auth/UPS
      • Fixed, but may go critical/unknown depending on timeout set up
      • Need check for passive, to define as down if checks not sent
    • SUSU/Boottime-SNMP
      • Needs to be rebooted to check
  • AOB
    • Contact us page needs new captcha
    • ECS moving sown.org.uk to new server soon
    • Site already made more secure
    • VOIP can be brought back.
      • need to fit 2U bracket to VOIP card
Facts about "2011/05/12"
Has date18:00:00, 12 May 2011 +
Has end date20:00:00, 12 May 2011 +
Has location59/1257 +